two planes started at the same time from the same airport and flew in oppposite directions. one flew 60 miles per hour faster than the other. in 5 hours, they were 2800 miles apart. find the rate of each plane.
Answer by checkley77(12844)

5X+5(X+60)=2,800
5X+5X+300=2,800
10X=2,800-300
10X=2,500
X=2,500/10
X=250 MPH FOR THE SLOWER PLANE.
250=60=310 MPH FOR THE FASTER PLANE.
PROOF:
5*250+5*310=2,800
1,250+1,550=2,800
2,800=2,800